OpenStudy (anonymous):

Graph the polar equation. r = 2 + 2sin θ Please help!

ganeshie8 (ganeshie8):

start by plotting few easy points like : \(\theta = 0, \pi, \pi/2, 3\pi/2\)

OpenStudy (anonymous):

Thank you so much!

ganeshie8 (ganeshie8):

u wlc :)